OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? a) 4x - y = 1 b) 4x + y = 1 c) x - 4y = -13 d)4x - y = -7

OpenStudy (alexandervonhumboldt2):

ok so you know slope formula?

OpenStudy (anonymous):

y1-y2 over x1-x2

OpenStudy (anonymous):

Well first find the slope, using the following formula: \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\] where your coordinates represent \[(x_1,y_1)~~\text{and}~~(x_2,y_2)\] Once you manage to get that, plug in m and your first coordinates in the point - slope form: \[y-y _{1}=m(x-x _{1})\] doing some algebra you can put it as the way it's provided in your answers.

OpenStudy (anonymous):

i got 4

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

oh wait nvm